Dorchester Company, on March 1, 2019 has a beginning Work in Process inventory of zero. All materials are added into production at the beginning of its production. There is only one production WIP inventory. On March 1, Dorchester started into production 17,500 units. At the end of the month there were 12,000 units completed and transferred into the Finished Goods Inventory. The ending WIP was 60% complete with respect to conversion. For the month of March the following costs were incurred and recorded in the WIP:

            Direct Material                      $26,000

            Direct Labor 27,000

            Factory Overhead 10,000

Dorchester uses the weighted-average process costing method. Use this information to determine the cost per equivalent unit of conversion for the month of March: (Round & enter final answers to the nearest cent.)
Answer #1

The cost per equivalent unit of conversion= $2.42

Working

Reconciliation of Units
A Beginning WIP                            -  
B Introduced                   17,500
C=A+B TOTAL                   17,500
D Transferred out                   12,000
E=C-D Ending WIP                     5,500

.

Statement of Equivalent Units(Weighted average)
Conversion
Units Complete % Equivalent units
Transferred out                                12,000 100%                   12,000
Ending WIP                                  5,500 60%                     3,300
Total                                17,500 Total                   15,300

.

Cost per Equivalent Units (Weighted average)
COST Conversion
Beginning WIP Inventory Cost (A) $ 0
Cost incurred during period (B) $ 37,000
Total Cost to be accounted for (C=A+B) $ 37,000
Total Equivalent Units(D)                     15,300
Cost per Equivalent Units (E=C/D) $ 2.42